# AdaR E T A K E : Adaptive Redundancy Reduction to Perceive Longer for Video-language Understanding

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) have revolutionized video understanding, yet are still limited by context length when processing long videos. Recent methods compress videos by leveraging visual redundancy uniformly, yielding promising results. Nevertheless, our quantitative analysis shows that redundancy varies significantly across time and model layers, necessitating a more flexible compression strategy. We propose AdaR E T A K E , a training-free method that flexibly reduces visual redundancy by allocating compression ratios among time and layers with theoretical guarantees. Integrated into state-of-the-art MLLMs, AdaR E T A K E improves processing capacity from 256 to 2048 frames while preserving critical information. Experiments on VideoMME, MLVU, LongVideoBench, and LVBench datasets demonstrate that AdaR E T A K E outperforms existing methods by 2.3% and 2.8% for 7B and 72B models, respectively, with even greater improvements of 5.9% and 6.0% on the longest LVBench. Our code is available at https://github.com/SCZwangxiao/video-FlexReduc.git.
@@ 1 Introduction
In pursuit of general intelligence, Multimodal Large Language Models (MLLMs) Li et al. (2024b); Lin et al. (2024) have revolutionized video understanding. However, current MLLMs require hundreds of tokens to represent a single image Wang et al. (2024a); Li et al. (2024a), limiting video lengths to less than 10 minutes Shen et al. (2024).
@@ 2.1 MLLM for Long Videos
Early attempts developed video agent systems Zhang et al. (2024a); Wang et al. (2024d); Luo et al. (2024) that segment videos into shorter clips and use MLLMs with open-source tools for retrieval, aggregation, and interpretation. However, a single model’s capabilities remain limited, reducing overall effectiveness. Length extrapolation methods Zhang et al. (2024c); Shang et al. (2024); Wei and Chen (2024) extend context windows beyond training lengths, but GPU memory still limits context size. To address this, Xue et al. introduced LongVILA, a multi-modal sequence parallelism system that distributes computation across GPUs, but this adds communication overhead Li et al. (2023b), affecting efficiency. In contrast, compression-based methods condense video tokens into shorter sequences. Approaches He et al. (2024a); Fei et al. (2024); Cheng et al. (2024); Zeng et al. (2024a); Man et al. (2024); Han et al. (2024) use Q-Former Li et al. (2023a) for token compression, reducing redundancy by leveraging language or query tokens. However, Q-Former, trained from scratch, lacks the world knowledge embedded in LLMs, making these methods suboptimal. Recent advances Shu et al. (2024); Shen et al. (2024); Liu et al. (2024); Wang et al. (2024c) integrate compression within the LLM, achieving promising results.
@@ 4.2 General Pipeline
Denote $T$ number of frames, $N$ number of tokens in each frame, $\tau$ number of frames in a chunk (can divide $T$), $L$ prompt length, and $C_{max}$ is a refined maximal context length.
Given raw frames and a text prompt as input, the visual encoder and projection layer derive video features $\mathbf{M}\in\mathbb{R}^{T\times N\times d}$, and the word embedding layer derives prompt features $\mathbf{P}\in\mathbb{R}^{L\times d}$. We split visual features into chunks of $\tau$ frames:
@@
Due to the autoregressive nature of LLMs, chunked prefilling is applied to each chunk, which is functionally equivalent to standard prefilling Zeng et al. (2024b). During the $i$-th iteration, chunk $i$ is first prefilled. For each layer $l$, the query states of the prompt $\mathbf{Q}^{(l)}_{i}\in\mathbb{R}^{L\times d}$ and the KV caches of chunk $i$ $\mathbf{K}^{(l)}_{i},\mathbf{V}^{(l)}_{i}\in\mathbb{R}^{h\times\tau N\times d}$ are stored, where $h$ is the number of heads. These, along with the chunk’s compression ratio $\alpha_{i}$, are processed by the layer-adaptive allocation module to determine the compression ratio for each layer:
@@ 4.3 Temporal-adaptive Allocation
$\alpha_{i}=\frac{C_{max}-L}{TN}\cdot\frac{\bar{d_{i}}}{\sum_{i=1}^{T/\tau}\bar{d_{i}}}.$ (9)
@@ 4.4 Layer-adaptive Allocation
In the $l$-th layer, we first calculate the attention score between prompt and the video tokens $\textbf{A}_{i}^{(l)}\in\mathbb{R}^{h\times L\times\tau N}$. We then calculate the head-averaged accumulated scores along all prompt tokens to measure the significance score of each token to the prompt, $\mathbf{a}_{i}^{(l)}\in\mathbb{R}^{\tau N}$:
$\mathbf{a}_{i}^{(l)}=\sum_{j=1}^{L}\frac{1}{h}\sum_{i=1}^{h}{\textbf{A}_{i}^{(l)}[i,j]}.$ (10)
@@ 5.1 Benchmarks and Implementations
Implementation Details. We integrated AdaR E T A K E into various MLLMs, including LLaVA-Video-7B Zhang et al. (2024e), QWen2VL-7B Wang et al. (2024a), QWen2.5VL-7B, and QWen2.5VL-72B. We densely sampled the video at 2 frames per second (fps), with a maximum of 2048 and 1024 frames for 7B and 72B models, respectively. For our main results (Section 5.2), we chose the maximal context length $C_{max}$ as 16K. In the ablation studies (Section5.3), we reduced the maximum number of sampled frames to 1024 and the context length to 1K without specification. The evaluation is conducted using LMMs-Eval Zhang et al. (2024b). All experiments were conducted on 8 Ascend 910B NPUs.
@@ 5.2 Main Results
4 +temporal allocation 1024 1K 55.1 65.6 44.8 +1.1
5 +scale up context length 1024 16K 56.0 71.7 48.0 +3.6
@@ 5.3 Ablation Studies
Ablation studies on temporal and layer-wise adaptive allocation. To identify the sources of performance improvements in our model, we conducted ablation studies, as summarized in Table 3. In the table, #0 represents the baseline model. In #1, we directly incorporates token compression into baseline model, and in #2, we increase the number of sampled frames while keeping the maximum context length fixed. In #3 and #4 we applly varying compression ratios across different layers and different frames respectively. Finally, #5 extends the context length. First, comparing rows 0,1 and 1,2 reveals that token compression introduces a slight performance drop (-1.1% on average). However, it enables the model to process more frames within the same context length, capturing richer information and ultimately yielding a net performance gain (2.5% on average versus -1.1%). 2) Comparing rows 2,3 and 3,4 shows that our strategy of distributing the compression ratio across frames and layers enhances performance (by 1.1% and 0.8% on average, respectively), confirming the effectiveness of our AdaR E T A K E . 3) Comparing rows 4 and 5 demonstrates that scaling the context length to the typical upper limit of MLLMs Shen et al. (2024) further improves performance significantly, with an average gain of 3.6%.
@@ A.1 Preliminaries
LLMs are characterized by an autoregressive generation mode, where each step involves using the last token to predict the next token. For a formal representation of our approach, we denote $\mathbf{K}^{(l)}$,$\mathbf{V}^{(l)}\in\mathbb{R}^{n-1\times d},i=1,2,...,L$ as the video KV cache in and $\mathbf{x}^{(l)}\in\mathbb{R}^{d}$ as the last token used as input at the current time step, where $n-1$ is the prefilled sequence length, $L$ is the number of layers and $d$ is the hidden state dimension. Note that we ignore the number of attention heads for simplification.
@@
where $\mathbf{W}_{Q}^{(l)},\mathbf{W}_{K}^{(l)},\mathbf{W}_{V}^{(l)}$ are transformation matrices. Then, the previous KV Cache is updated:

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have revolutionized video understanding, yet are still limited by context length when processing long videos. Recent methods compress videos by leveraging visual redundancy uniformly, yielding promising results. Nevertheless, our quantitative analysis shows that redundancy varies significantly across time and model layers, necessitating a more flexible compression strategy. We propose AdaR E T A K E , a training-free method that flexibly reduces visual redundancy by allocating compression ratios among time and layers with theoretical guarantees. AdaR E T A K E can be seamlessly integrated into existing MLLMs as a plug-and-play solution, extending their processing capacity from 256 to 2048 frames while preserving critical information. Experiments on VideoMME, MLVU, LongVideoBench, and LVBench datasets demonstrate that AdaR E T A K E outperforms existing methods by 2.3% and 2.8% for 7B and 72B models, respectively, with even greater improvements of 5.9% and 6.0% on the longest LVBench. Our code is available at https://github.com/SCZwangxiao/video-FlexReduc.git.
@@ 1 Introduction
In pursuit of general intelligence, Multimodal Large Language Models (MLLMs) Li et al. (2024b); Lin et al. (2024); Wang et al. (2025b, a) have revolutionized video understanding. However, current MLLMs require hundreds of tokens to represent a single image Wang et al. (2024a); Li et al. (2024a); Wang et al. (2023), limiting video lengths to less than 10 minutes Shen et al. (2024); Gan et al. (2023).
@@ 2.1 MLLM for Long Videos
Early attempts developed video agent systems Zhang et al. (2024a); Wang et al. (2024d); Luo et al. (2024); Liu et al. (2018) that segment videos into shorter clips and use MLLMs with open-source tools for retrieval, aggregation, and interpretation. However, a single model’s capabilities remain limited, reducing overall effectiveness. Length extrapolation methods Zhang et al. (2024c); Shang et al. (2024); Wei and Chen (2024) extend context windows beyond training lengths, but GPU memory still limits context size. To address this, Xue et al. introduced LongVILA, a multi-modal sequence parallelism system that distributes computation across GPUs, but this adds communication overhead Li et al. (2023b), affecting efficiency. In contrast, compression-based methods condense video tokens into shorter sequences. Approaches He et al. (2024a); Fei et al. (2024); Cheng et al. (2024); Zeng et al. (2024a); Man et al. (2024); Han et al. (2024) use Q-Former Li et al. (2023a) for token compression, reducing redundancy by leveraging language or query tokens. However, Q-Former, trained from scratch, lacks the world knowledge embedded in LLMs, making these methods suboptimal. Recent advances Shu et al. (2024); Shen et al. (2024); Liu et al. (2024); Wang et al. (2024c) integrate compression within the LLM, achieving promising results.
@@ 4.2 General Pipeline
Denote $T$ number of frames, $N$ number of tokens in each frame, $\tau$ number of frames in a chunk (can divide $T$), $S$ prompt length, $L$ number of LLM layers, and $C_{max}$ is a refined maximal context length.
Given raw frames and a text prompt as input, the visual encoder and projection layer derive video features $\mathbf{M}\in\mathbb{R}^{T\times N\times d}$, and the word embedding layer derives prompt features $\mathbf{P}\in\mathbb{R}^{S\times d}$. We split visual features into chunks of $\tau$ frames:
@@
Due to the autoregressive nature of LLMs, chunked prefilling is applied to each chunk, which is functionally equivalent to standard prefilling Zeng et al. (2024b). During the $i$-th iteration, chunk $i$ is first prefilled. For each layer $l$, the query states of the prompt $\mathbf{Q}^{(l)}_{i}\in\mathbb{R}^{S\times d}$ and the KV caches of chunk $i$ $\mathbf{K}^{(l)}_{i},\mathbf{V}^{(l)}_{i}\in\mathbb{R}^{h\times\tau N\times d}$ are stored, where $h$ is the number of heads. These, along with the chunk’s compression ratio $\alpha_{i}$, are processed by the layer-adaptive allocation module to determine the compression ratio for each layer:
@@ 4.3 Temporal-adaptive Allocation
$\alpha_{i}=\frac{C_{max}-S}{TN}\cdot\frac{\bar{d_{i}}}{\sum_{i=1}^{T/\tau}\bar{d_{i}}}.$ (9)
@@ 4.4 Layer-adaptive Allocation
In the $l$-th layer, we first calculate the attention score between prompt and the video tokens $\textbf{A}_{i}^{(l)}\in\mathbb{R}^{h\times S\times\tau N}$. We then calculate the head-averaged accumulated scores along all prompt tokens to measure the significance score of each token to the prompt, $\mathbf{a}_{i}^{(l)}\in\mathbb{R}^{\tau N}$:
$\mathbf{a}_{i}^{(l)}=\sum_{j=1}^{S}\frac{1}{h}\sum_{i=1}^{h}{\textbf{A}_{i}^{(l)}[i,j]}.$ (10)
@@ 5.1 Benchmarks and Implementations
Implementation Details. We integrated AdaR E T A K E into various MLLMs, including LLaVA-Video-7B Zhang et al. (2024e), QWen2VL-7B Wang et al. (2024a), QWen2.5VL-7B, and QWen2.5VL-72B. We densely sampled the video at 2 frames per second (fps), with a maximum of 2048 and 1024 frames for 7B and 72B models, respectively. For our main results (Section 5.2), we chose the maximal context length $C_{max}$ as 16K. In the ablation studies (Section5.3), we reduced the maximum number of sampled frames to 1024 and the context length to 1K without specification. The evaluation is conducted using LMMs-Eval Zhang et al. (2024b).
@@ 5.2 Main Results
4 +temporal allocation 1024 1K 55.1 65.6 44.8 +1.0
5 +scale up context length 1024 16K 56.0 71.7 48.0 +3.4
6 +scale up frames 2048 16K 56.4 72.0 48.9 +0.6
@@ 5.3 Ablation Studies
Ablation studies on temporal and layer-wise adaptive allocation. To identify the sources of performance improvements in our model, we conducted ablation studies, as summarized in Table 3. In the table, #0 represents the baseline model. In #1, we directly incorporates token compression into baseline model, and in #2, we increase the number of sampled frames while keeping the maximum context length fixed. In #3 and #4 we applly varying compression ratios across different layers and different frames respectively. Finally, #5 extends the context length. First, comparing rows 0,1 and 1,2 reveals that token compression introduces a slight performance drop (-0.8% on average). However, it enables the model to process more frames within the same context length, capturing richer information and ultimately yielding a net performance gain (2.5% on average versus -0.8%). 2) Comparing rows 2,3 and 3,4 shows that our strategy of distributing the compression ratio across frames and layers enhances performance (by 1.0% and 0.8% on average, respectively), confirming the effectiveness of our AdaR E T A K E . 3) Comparing rows 4 and 5 demonstrates that scaling the context length to the typical upper limit of MLLMs Shen et al. (2024) further improves performance significantly, with an average gain of 3.4%.
@@ 7 Limitations
8 Acknowledgment
This work was supported in part by the National Natural Science Foundation of China(Grant Nos. 62376069 and 62236003), in part by the Young Elite Scientists Sponsorship Program by CAST (Grant No. 2023QNRC001), in part by Guangdong Basic and Applied Basic Research Foundation (Grant No. 2024A1515012027), in part by Jiangsu Science and Technology Major Program (Grant No. BG2024041), and in part by the Shenzhen Science and Technology Program (Grant Nos. KQTD20240729102207002 and ZDSYS20230626091203008).
@@ A.1 Preliminaries
LLMs are characterized by an autoregressive generation mode, where each step involves using the last token to predict the next token. For a formal representation of our approach, we denote $\mathbf{K}^{(l)}$,$\mathbf{V}^{(l)}\in\mathbb{R}^{(n-1)\times d},i=1,2,...,L$ as the video KV cache in and $\mathbf{x}^{(l)}\in\mathbb{R}^{d}$ as the last token used as input at the current time step, where $n-1$ is the prefilled sequence length, $L$ is the number of layers and $d$ is the hidden state dimension. Note that we ignore the number of attention heads for simplification.
@@
where $\mathbf{W}_{Q}^{(l)},\mathbf{W}_{K}^{(l)},\mathbf{W}_{V}^{(l)}\in\mathbb{R}^{d\times d}$ are transformation matrices. Then, the previous KV Cache is updated: